# Enhancing the Robustness of LLM-Generated Code: Empirical Study and Framework

### v3

@@ 4.2. RQ6: Efficiency
Greedy Sampling 1.52 3.03 1.83 1.13 3.01
RobGen-Adj 1.53 (+0.6%) 3.05 (+0.7%) 1.83 (+0.0%) 1.18 (4.4%) 3.02 (0.3%)
RobGen-Ins 2.72 (+78.9%) 4.21 (+38.9%) 2.09 (+14.2%) 1.77 (+56.6%) 3.99 (+32.6%)
RobGen-Adj+Ins 1.99 (+30.9%) 3.89 (+28.3%) 1.96 (+7.1%) 1.38 (+22.1%) 3.30 (+9.6%)
@@ 4.2.2. Results
The execution time results are presented in Table4. Among all methods, RobGen-Ins exhibits the highest time overhead. For instance, after applying RobGen-Ins, DeepSeekCoder-1.3B ’s execution time increased from 1.52 to 2.72, representing a 78.9% increase compared to Greedy Sampling. Similarly, for StarCoder2-7B, execution time increased by 32.6% after using RobGen-Ins. This trade-off is considered acceptable, as RobGen-Ins involves an additional model invocation to generate the missing content. In contrast, RobGen-Adj has minimal impact on execution time. For example, when using RobGen-Adj, DeepSeekCoder-6.7B experienced only a 4.4% increase, while Qwen2.5-Coder-1.5B saw a negligible 0.6% increase. Furthermore, RobGen-Adj+Ins results in lower time overhead compared to using RobGen-Ins alone. For instance, when DeepSeekCoder-1.3B employed RobGen-Adj+Ins, the additional time overhead was reduced from 78.9% to 30.9%, demonstrating that RobGen-Adj effectively preemptively addresses robustness issues, thereby reducing the need for additional insertions during RobGen-Ins.
@@ 5. Threats to validity
Internally, a potential threat is the use of LLMs in our research. To mitigate biases, we use officially released models, deploy them per publishers’ guidelines, and validate outputs through tests. We also ensure fairness by using consistent prompts and parameters. Another concern is the representativeness and quality of code data. To address this, we use the CoderEval benchmark for realistic code generation. Due to resource constraints, we evaluated only four models and focused on Java, limiting generalizability. More extensive studies are needed. While RobGen was designed based on findings from these four models, we included Starcoder in experiments to demonstrate broader applicability. Externally, we minimize human bias in robustness analysis by using quantifiable metrics instead of manual judgment.
@@ 6.1. LLM-based Code Generation
LLMs have shown remarkable performance in code generation, with applications in automated code generation (Yuan et al., 2023; Ugare et al., 2024; Su and McMillan, 2024; Liu et al., 2024a; Wang et al., 2024c), code translation (Yan et al., 2023; Yang et al., 2024a; Ou et al., 2025; Wang et al., 2024b), commit message generation (Tao et al., 2021; He et al., 2023), unit test generation (Schäfer et al., 2024; Lops et al., 2024; Yuan et al., 2024), and defect localization (An et al., 2024; Ji et al., 2024; Du et al., 2024b). Code-specific LLMs, like DeepSeekCoder (Guo et al., 2024c), StarCoder2 (Lozhkov et al., 2024), and Qwen2.5Coder (Yang et al., 2024b), excel in code-related tasks, with high-quality benchmarks such as HumanEval (Chen et al., 2021b), MBPP (Austin et al., 2021), Classeval (Du et al., 2024a) and CodeEval (Yu et al., 2023) assessing their capabilities. While most studies focus on improving code generation and evaluating quality with the Pass@k metric (Chen et al., 2021b), some explore code security (Li et al., 2024; Islam et al., 2024b), robustness(Sarker et al., 2024; Awal et al., 2024), hallucination(Zhang et al., 2025) and trustworthy(Wang et al., 2024a). However, existing benchmarks fall short in assessing the robustness of model-generated code. This work compares the robustness of LLM-generated code with human-written code, identifying common deficiencies and offering insights for future improvements.
@@ 6.2. Robustness of LLM-generated Code
Previous work lacks in-depth analysis of robustness issues in LLM-generated code, particularly in real-world tasks. Many methods are either impractical or involve high overhead. In contrast, our study provides a deeper empirical analysis and identifies common robustness issues across multiple models. We propose RobGen, a model-agnostic plug-in framework that improves robustness with minimal overhead, enhancing code quality during both decoding and post-generation in a lightweight, adaptable manner.
@@ 7. CONCLUSION
In conclusion, this study emphasizes the need for robustness in LLM-generated code, which is often overlooked in favor of correctness. Using the CoderEval benchmark, we identify critical gaps, such as missing null checks. To address these, we propose RobGen, a plug-in framework that enhances code robustness without retraining, through techniques like RobGen-Adj and RobGen-Ins. Our experiments show a 20% reduction in less robust code, demonstrating the effectiveness of this approach. RobGen offers a flexible, model-agnostic solution for improving LLM-generated code reliability across various tasks.

### v4

@@ 4.2. RQ6: Efficiency
Greedy Sampling 4.67 14.09 2.41 1.41 14.79
RobGen w/o Checker 4.78(+2.3%) 14.16(+0.4%) 2.74(+13.7%) 1.45(+1.3%) 14.83(+0.2%)
RobGen 6.23(+33.4%) 15.77(+11.4%) 7.18(+197.9%) 1.65(+17.0%) 16.50(+11.6%)
@@ 4.2.2. Results
The execution time results are presented in Table4. Among all methods, RobGen exhibits the highest time overhead. For instance, after applying RobGen, DeepSeekCoder-1.3B ’s execution time increased from 4.67 to 6.23, representing a 33.4% increase compared to Greedy Sampling.
In contrast, RobGen w/o Checker has minimal impact on execution time. For example, when using RobGen w/o Checker, DeepSeekCoder-6.7B experienced only a 0.4% increase, while StarCoder2-7B saw a negligible 0.2% increase. Although this trade-off is acceptable given that RobGen requires an additional model invocation, we note that applying RobGen increases the time overhead of Qwen2.5-Coder-1.5B to nearly twice that of Greedy Sampling. In contrast, RobGen w/o Checker introduces minimal overhead. This is because the checker operates at the line level, making it incompatible with KV cache acceleration(Yang et al., 2025a) and thus significantly increasing inference time for longer code. Excessive code generation by Qwen2.5-Coder-1.5B leads to longer preceding code, which in turn increases the Checker’s inference time.
Finding 6: RobGen adds only a modest time overhead (+33.4%), highlighting its efficiency, whereas the Checker’s overhead can rise sharply (+197.9%) due to model over-generation (repetition), occasionally resulting in abnormal behavior.
@@ 5. THREATS TO VALIDITY
Internally, a potential threat is the use of LLMs in our research. To mitigate biases, we use officially released models, deploy them per publishers’ guidelines, and validate outputs through tests. We also ensure fairness by using consistent prompts and parameters. Another concern is the representativeness and quality of code data. To address this, we use the CoderEval benchmark for realistic code generation. Due to resource constraints, we evaluated only four models and focused on Java, limiting generalizability. More extensive studies are needed. While RobGen was designed based on findings from these four models, we included Starcoder in experiments to demonstrate broader applicability. To minimize human bias in robustness evaluation, we use an LLM Evaluator instead of manual judgment. Each task is evaluated three times with randomized input order of human-written and model-generated code to reduce LLM-induced randomness. We tested two prompt designs—pairwise comparison and single-score rating—and found pairwise comparison best aligned with human judgments, which we adopted as our primary evaluation method.
@@ 6.1. LLM-based Code Generation
LLMs have shown remarkable performance in code generation, with applications in automated code generation (Yuan et al., 2023; Ugare et al., 2024; Su and McMillan, 2024; Liu et al., 2024a; Wang et al., 2024c), code translation (Yan et al., 2023; Yang et al., 2024a; Ou et al., 2025; Wang et al., 2024b), commit message generation (Tao et al., 2021; He et al., 2023), unit test generation (Schäfer et al., 2024; Lops et al., 2024; Yuan et al., 2024), and defect localization (An et al., 2024; Ji et al., 2024; Du et al., 2024b). Code-specific LLMs, like DeepSeekCoder (Guo et al., 2024c), StarCoder2 (Lozhkov et al., 2024), and Qwen2.5Coder (Yang et al., 2024b), excel in code-related tasks, with high-quality benchmarks such as HumanEval (Chen et al., 2021b), MBPP (Austin et al., 2021), Classeval (Du et al., 2024a) and CodeEval (Yu et al., 2023) assessing their capabilities. While most studies focus on improving code generation and evaluating quality with the Pass@k metric (Chen et al., 2021b), some explore code security (Li et al., 2024a; Islam et al., 2024b), robustness(Sarker et al., 2024; Awal et al., 2024), hallucination(Zhang et al., 2025) and trustworthy(Wang et al., 2024a). However, existing benchmarks fall short in assessing the robustness of model-generated code. This work compares the robustness of LLM-generated code with human-written code, identifying common deficiencies and offering insights for future improvements.
@@ 6.2. Robustness of LLM-generated Code
Previous work lacks in-depth analysis of robustness issues in LLM-generated code, particularly in real-world tasks. Many methods are either impractical or involve high overhead. In contrast, our study provides a deeper empirical analysis and identifies common robustness issues across multiple models. We propose RobGen, a model-agnostic plug-in framework that improves robustness with minimal overhead, enhancing code quality during decoding in a lightweight, adaptable manner.
@@ 7. CONCLUSION
In conclusion, this study emphasizes the need for robustness in LLM-generated code, which is often overlooked in favor of correctness. Using the CoderEval benchmark, we identify critical gaps, such as missing null checks. To address these, we propose RobGen, a plug-in framework that enhances code robustness without retraining, through techniques like Decoding Adjustment. Our experiments show a 10% reduction in less robust code, demonstrating the effectiveness of this approach. RobGen offers a flexible, model-agnostic solution for improving LLM-generated code reliability across various tasks.